OpenStudy (3psilon):

Dad dreamed that he put $1 on the first square,$2 on the second, $4 dollars on the third, and so on, doubling the amount each time.If it costs $65,535 to cover all the squares,how many squares were in Dad's dream?

OpenStudy (anonymous):

\[1+2+2^2+2^3+...+2^{n-1}=2^n-65535\]

OpenStudy (anonymous):

oops i mean \[=2^n-1=65535\]

OpenStudy (3psilon):

My book says there are 16 squares but only 1 covered ?

OpenStudy (anonymous):

yes there are 16

OpenStudy (anonymous):

\[2^n-1=65535\] \[2^n=65526\] and \[2^{16}=65536\]

OpenStudy (anonymous):

another typo, \(2^n=65536\)

OpenStudy (anonymous):

It's a geometric series, right? So in general we use: \[ \Large \sum_{i=0}^{n-1}a_0r^i = a_0\frac{1-r^n}{1-r} \]

OpenStudy (anonymous):

Where \(a_0 = 1\) and \(r=2\)...

OpenStudy (anonymous):

65535 = 2^16 - 1 --> 16 squares